geerky42 (geerky42):

A body is acted upon by a force of 10 N and undergoes a displacement in the direction of the force in accordance with the relation s = 3t² + 2t, where s is the displacement in meters and t is in seconds. The rate at which the force is doing work at the instant t = 2s is... A) 14 W B) 160 W C) 120 W D) 140 W E) 12 W

OpenStudy (anonymous):

put t =2 in s=3(2)^2+2*2 s=16 now w=Fd cos 0 because force and displacement in same direction

OpenStudy (anonymous):

W=10*16 W=160

geerky42 (geerky42):

It's funny because unit is watt, so I think I was supposed to find power or something. W = 160 J...

OpenStudy (anonymous):

power is the rate at which work is done meaning it the workdone per time unit. one watt =1joule/second. work done=160j/s therefore power =160w

geerky42 (geerky42):

Why can't I use \(\int\limits \!\! \int\limits F \text{ d}\! s \text{ d}\! t \) where \(W = \int\limits F \text{ d}\! s\)

OpenStudy (anonymous):

There is no way around the fact that we have calculated the work done after 2 seconds as 160J. This is NOT the current power output. I think you need to differentiate the function, s(t) = 3t² + 2t to get the rate of displacement. This can give you the rate of work done (i.e. power). I.e\[Power = \frac{ W }{ dt } = F \frac{ s(t) }{ dt }\] The answer I get is one of the options..

geerky42 (geerky42):

So... the answer is 140 W? This is what I got.

geerky42 (geerky42):

@furnessj

geerky42 (geerky42):

@mayankdevnani @mahmit2012

OpenStudy (mayankdevnani):

answer is 160W

OpenStudy (mayankdevnani):

s = 3t² + 2t, t =2 in ___ s=3(2)^2+2*2

OpenStudy (mayankdevnani):

ok @geerky42

OpenStudy (mayankdevnani):

s=16

OpenStudy (mayankdevnani):

Work done=Force*Displacemnt W=F*s W=10N*16m W=160Nm=160J

OpenStudy (mayankdevnani):

you find that The rate at which the force is doing work at the instant t = 2s is... so it is POWER. at any instant convert J(Joule) into W(Watt) because question says at instant

OpenStudy (mayankdevnani):

ok @geerky42

geerky42 (geerky42):

I see. Thanks.

OpenStudy (mayankdevnani):

welcome

OpenStudy (anonymous):

I disagree with you mayankdevani, the units are simply not dimensinally correct for power using force x distance moved. Power is also defined as F x velocity, and ds/dt is the velocity... http://hyperphysics.phy-astr.gsu.edu/hbase/pow.html

geerky42 (geerky42):

@furnessj has good point. Can you defend yourself and explain and why you think your answer is correct? @mayankdevnani